OpenStudy (jazzyfa30):

convert each degree measure into radians and each radian measure into degrees. A)30 degrees B)-285 degrees C) 4 pi/3 D)975 degrees E)-10 degrees

OpenStudy (jazzyfa30):

@johnweldon1993

OpenStudy (johnweldon1993):

Alright: To convert from Radians to degrees \[\large \text{Multiply by} \frac{180}{\pi}\] To convert from Degrees to Radians \[\large \text{Multiply by} \frac{\pi}{180}\]

OpenStudy (johnweldon1993):

So lets look at examples A and C A) 30 degrees \[\large 30^\circ \times \frac{\pi}{180^\circ} = \frac{30^\cancel{\circ} \pi}{180^\cancel{\circ}} = \frac{\pi}{6}\] Right? Because simplifying the fraction 30/180 we get 1/6 so that would be pi/6

OpenStudy (johnweldon1993):

And now C C) 4pi/3 \[\large \frac{4\pi}{3} \times \frac{180^\circ}{\pi} = \frac{720^\circ\cancel{\pi}}{3\cancel{\pi}} = 240^\circ\]

OpenStudy (jazzyfa30):

ok so is that how i do B,D, and E like A

OpenStudy (johnweldon1993):

Correct

OpenStudy (jazzyfa30):

thanks